• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jwexler

    Log entry time 07:59:01 on June 5,2010

    Entry number 324650

    keyword=Shift Summary: Owl, 06/05/2010

    Owl Shift Summary - 06/05/2010

    00:00 - Beam still down from previous shift. Mysterious iochacp2 issue reported by MCC unresolved (previous HALOG entries suggest this is a Compton IOC).

    01:15 - MCC finished recovering beam. Beginning beam recovery on PbD#3

    01:55 - Heater off, cooler on, beam recovery procedure completed, ramping up to 50uA. Beginning production runs (after two runs w/ coda crashes)

    03:30 - LeCroy HV crates no longer responding - white connectivity alarm starts. Attempt fixing via rebooting IOC 11 and IOC 22.

    04:30 - Still no response from HV crates. Attempt second reboot

    05:00 - RC called, is looking into HV crate alarms remotely.

    05:30 - HV crates back online.

    06:00 - Change to slug 29, HWP OUT.

    Runs:
    4290-1: Coda crash
    4292-6: Production (IHWP IN, slug 28)
    4297-9:Production (IHWP OUT, slug 29)